### v1

# StackTrans: From Large Language Model to Large Pushdown Automata Model

Abstract
The Transformer architecture has emerged as a landmark advancement within the broad field of artificial intelligence, effectively catalyzing the advent of large language models (LLMs). However, despite its remarkable capabilities and the substantial progress it has facilitated, the Transformer architecture still has some limitations. One such intrinsic limitation is its inability to effectively capture the Chomsky hierarchy, such as regular expressions or deterministic context-free grammars. Drawing inspiration from pushdown automata, which efficiently resolve deterministic context-free grammars using stacks, we propose StackTrans to address the aforementioned issue within LLMs. Unlike previous approaches that modify the attention computation, StackTrans explicitly incorporates hidden state stacks between Transformer layers. This design maintains compatibility with existing frameworks like flash-attention. Specifically, our design features stack operations – such as pushing and popping hidden states – that are differentiable and can be learned in an end-to-end manner. Our comprehensive evaluation spans benchmarks for both Chomsky hierarchies and large-scale natural languages. Across these diverse tasks, StackTrans consistently outperforms standard Transformer models and other baselines. We have successfully scaled StackTrans up from 360M to 7B parameters. In particular, our from-scratch pretrained model StackTrans-360M outperforms several larger open-source LLMs with 2–3$\times$ more parameters, showcasing its superior efficiency and reasoning capability.
1 Introduction
In the era of large language models (LLMs) [GPT-4, 2023], the Transformer architecture has emerged as the nearly universal backbone, achieving remarkable success across various domains and beyond [Bi et al., 2024; Bai et al., 2023; Zhang et al., 2024a]. However, recent empirical studies [Delétang et al., 2022; Hahn, 2020] have demonstrated that Transformers struggle with tasks involving Chomsky hierarchies [Chomsky, 1956], such as regular expressions (REs) and deterministic context-free grammars (DCFs). For example, in an RE matching task, each training example consists of a string within a specified length range, paired with a matching-or-not label. While Transformers can perform well within the length boundaries of the training set, they often fail to maintain consistent performance when evaluated on input strings that are longer or shorter than those in the training data. One theoretical explanation is that standard Transformers lack inductive biases [Mitchell, 1980; Battaglia et al., 1806; Sartran et al., 2022]. Without inherent assumptions to guide learning, Transformers struggle to effectively capture the underlying grammar of the Chomsky hierarchy in the training set. Consequently, they cannot generalize beyond the training data and perform poorly on inputs of lengths that differ from those seen during training. On the other hand, natural languages are generally considered to belong to a class of languages that extends beyond DCF [Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985]. This inherent limitation of Transformers may possibly hinder their application in real-world natural language modeling. Such deficiencies may also have the risk to block LLMs to achieve more advanced forms of intelligence.
In the realm of formal language theory and computational linguistics [Chomsky, 1956; Savage and Computation, 1998; Sipser, 1996], it is widely recognized that automata augmented with stacks correspond to different levels of the Chomsky hierarchy grammars  [^1] The Chomsky grammars of type 3 (REs) and type 2 (DCFs) can be recognized by pushdown automata (automata equipped with a stack) [Chomsky, 1956; Chomsky and Schützenberger, 1959]. With proper enhancement, pushdown automata can even resolve type 1 (context-sensitive, CS) and type 0 (unrestricted) grammars – 2-stack automata (automata equipped with two independent stacks) are equivalent to Turing machines, and can recognize type 0 grammars [Yau, 1969].. Given the success of stack-equipped automata in handling rather complex grammars, it is a natural progression to incorporate the data structure of stack into the Transformer architecture. Recently, researchers have proposed the stacked attention mechanism [DuSell and Chiang, 2024; Li et al., 2024] and have examined its viability upon multiple relatively small benchmarks.
Drawing inspiration from pushdown automata and prior research, we introduce hidden state stacks into the Transformer architecture, proposing a novel method, StackTrans. Unlike the stacked attention mechanisms [DuSell and Chiang, 2024; Li et al., 2024], which replace the standard attention, StackTrans incorporates differentiable stacks between the Transformer layers, meanwhile preserving the integrity of the Transformer layers (Figure LABEL:fig:method:a). This integration allows us to embed the assumptions of the Chomsky hierarchy into the model, enabling StackTrans to inherently model and learn REs and DCFs. Moreover, this design maintains compatibility with existing frameworks like flash-attention [Dao et al., 2022; Dao, 2024], enabling seamless integration into efficient LLM training pipelines. Specifically, the stack stores hidden states generated by the preceding layer and updates through operations such as push and pop at each decoding time step. To enable end-to-end training of StackTrans, we design soft stack operations, thereby making the hidden state stack differentiable. StackTrans also adopts a multi-head stack to improve its representation capability. Additionally, we find that the standard stack reading operation (which only returns the top element in the stack) may cause unstable training. Therefore, we propose the global reading operation through a learnable query-over-stack attention, stabilizing the training process and enriching the expressiveness of StackTrans.
We conduct comprehensive experiments on multiple benchmarks spanning both formal languages [Delétang et al., 2022] and natural languages [Groeneveld et al., 2024]. Specifically, in RE and DCF tasks, StackTrans outperforms the standard Transformer by at least 30%, achieving nearly 100% test accuracy in most scenarios, as shown in Figure LABEL:fig:method:b. Surprisingly, during natural language evaluation, StackTrans also demonstrates substantial improvements on tasks such as common sense reasoning and question answering. Furthermore, we have successfully scaled StackTrans up from 360M to 7B parameters. In particular, our open-sourced StackTrans-360M, which is pretrained on a corpus of approximately 1T tokens, performs better than or comparably to state-of-the-art LLMs with 2–3$\times$ more parameters, as shown in Figure LABEL:fig:method:c.
The contribution of this paper can be summarized as below:
  • We introduce StackTrans, which incorporates hidden state stacks between Transformer layers. This integration enables StackTrans to inherently learn grammars from the Chomsky hierarchy, including REs and DCFs.
  • We design the hidden state stack to be differentiable, which employs soft push, pop, and no-op operations, a multi-head stack mechanism, and a global stack reading operation. These innovations ensure stable and end-to-end training of StackTrans.
  • We conduct extensive experiments on multiple formal language benchmarks, demonstrating StackTrans’s effectiveness in learning Chomsky hierarchy grammars such as REs and DCFs. Furthermore, we have successfully scaled StackTrans up from 360M to 7B parameters for general language modeling. Evaluations on large-scale natural language benchmarks show that StackTrans-360M outperforms baselines with even 23×2-3$\times$2 - 3 × more parameters.
2 Background
The foundational success of LLMs can be attributed to the development of the Transformer architecture [Vaswani et al., 2017] and its numerous variations, which serve as the backbone for LLMs [Radford et al., 2018; GPT-4, 2023; GPT-4o, 2024]. Despite their widespread adoption, the Transformer architecture has inherent expressivity limitations [Hahn, 2020]. Recent studies have shown that standard Transformers struggle with recursive and hierarchical patterns across both synthetic and real-world tasks [Joulin and Mikolov, 2015; Grefenstette et al., 2015; Sartran et al., 2022], and need to equip neural networks with external data structures. These limitations pose a potential risk to the natural language modeling capabilities of Transformers, as suggested by discussions surrounding the classification of natural languages within the Chomsky hierarchy [Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985]. For more detailed related work, please refer to §A. Based on the same principle, StackTrans introduces differentiable hidden state stacks in a modular and scalable manner. Without altering the attention mechanism, StackTrans integrates stack operations into layer-wise hidden state updates. This design maintains compatibility with existing frameworks like flash-attention [Dao et al., 2022; Dao, 2024] and supports architectures ranging from 0.36B to 7B parameters. By learning stack operations explicitly, StackTrans is capable of addressing broader linguistic and algorithmic challenges.
Figure 3: Illustration of the multi-head differentiable stack. Our designed differentiable stack includes the updating mechanism based on three action probabilities (i.e., push, pop, and no-op), the stack mask maintenance, as well as a gated global reading mechanism. To improve both memory efficiency and representational flexibility, we also add multi-head and low-rank mechanisms.
3 Method
As introduced previously, the standard Transformer architecture [Vaswani et al., 2017] struggles to learn the Chomsky hierarchy due to its lack of inductive biases. To address this limitation, we propose StackTrans, which incorporates hidden state stacks into the Transformer architecture. These stacks introduce the assumptions of the Chomsky hierarchy, enhancing the model’s ability to capture hierarchical structures. In StackTrans, the hidden state stacks augment the information flow by routing token-level hidden states through learnable stack operations, such as stack updating and reading. Specifically, the stacks are integrated between standard Transformer layers, where they store hidden states and perform updates and readings via soft operations (see §3.1 for details). To further improve the expressiveness of StackTrans, we introduce a multi-head stack mechanism (please refer to §3.2). This design enhances the ability of StackTrans to capture diverse patterns with low-rank representations. Finally, to ensure robust training and avoid issues such as stack operation collapsing, we introduce stack regularization techniques. We also implement stack truncation to facilitate batching and parallel training (§3.3). These innovations and designs collectively enhance the effectiveness and stability of StackTrans in learning Chomsky hierarchy grammars.
3.1 Hidden State Stack
In computational linguistics, it is well established that the Chomsky hierarchy grammars can be resolved by different classes of automata, with pushdown automata being the minimal computational model for DCFs [Chomsky and Schützenberger, 1959]. Drawing inspiration from this, we incorporate stacks into the Transformer architecture to address its limitations in learning the Chomsky hierarchy. Recall that a standard stack is a last-in-first-out storage structure that allows the top element to be read or updated (i.e., operations like peek, push, and pop). Given a hidden state sequence $h_{1},\cdots,h_{l}$ (for now, we do not consider any computational dependencies among $h_{t}$’s), at each time-step $t$, StackTrans is designed to either push the current hidden state $h_{t}$ into the stack or pop the top element from the stack, and then peek the current top element (at this stage, we ignore how these operations are determined or executed by the model). This procedure results in a new stack-operated sequence, which is a permutation of $h_{1},\cdots,h_{l}$. Ideally, if the stack operations are correct, it is highly plausible that the model can effectively learn the target Chomsky hierarchy grammar.
Since the hidden state stack will be incorporated into StackTrans, it must be differentiable to enable end-to-end training. However, standard stack operations such as push and pop are discrete, which disrupts gradient back-propagation and hinders the training process. To address this challenge, we introduce a soft operation mechanism, following earlier explorations [Grefenstette et al., 2015; Joulin and Mikolov, 2015]. In this mechanism, the results of the stack operations are continuously interpolated based on some trainable parameters. This design not only makes the operations differentiable but also allows the model to learn the stack operations through these trainable parameters.
Soft update
The stack at time-step $t$ can be formalized as a list of vectors, where $\text{St}_{t}[i]$ refers to the $i$-th element from top to bottom in the current stack $\text{St}_{t}$. We define three candidate operations for StackTrans determined by the current hidden state $h_{t}$ (here, $\text{St}_{t}[i]$ and $h_{t}$ both belong to $\mathbb{R}^{d}$, meaning they share the same width $d$): ❶ “push” shifts every element down by one position and puts $h_{t}$ at the top; ❷ “pop” removes the top element and moves every remaining element up by one position; and ❸ “no-op” does not alter $\text{St}_{t}$ at all. Instead of discretely selecting one of these operations, the soft update mechanism computes a distribution over the candidate operations. This is achieved through a learned linear projection $A\in\mathbb{R}^{3\times d}$ followed by a softmax function:
$a_{t}=[a_{t}^{\text{push}};a_{t}^{\text{pop}};a_{t}^{\text{noop}}]=\text{Softmax}(Ah_{t})$ (1)
where each scalar in $a_{t}$ represents the probability corresponding to one operation. We then combine the results of each operation based on their respective probabilities to update the stack as follows:
$\text{St}_{t+1}[i]=\left\{\begin{array}[]{lclcll}a_{t}^{\text{push}}\cdot h_{t}&+&a_{t}^{\text{pop}}\cdot\text{St}_{t}[1]&+&a_{t}^{\text{noop}}\cdot\text{St}_{t}[0],&\text{if }i=0\\[3.0pt] a_{t}^{\text{push}}\cdot\text{St}_{t}[i-1]&+&a_{t}^{\text{pop}}\cdot\vec{0}&+&a_{t}^{\text{noop}}\cdot\text{St}_{t}[i],&\text{if }i=S-1\\[3.0pt] a_{t}^{\text{push}}\cdot\text{St}_{t}[i-1]&+&a_{t}^{\text{pop}}\cdot\text{St}_{t}[i+1]&+&a_{t}^{\text{noop}}\cdot\text{St}_{t}[i],&\text{otherwise}\\[3.0pt] \end{array}\right.$ (2)
where $S$ denotes the size of the stack. The first and the second rows in Equation 2 correspond to the top element ($\text{St}_{t+1}[0]$) and the bottom element ($\text{St}_{t+1}[S-1]$) respectively, while the last row pertains to the intermediate elements. It is important to note that a zero vector is always maintained at the bottom of the stack, as indicated in the “pop” term of the second row in Equation 2.
The soft update mechanism is fully differentiable, enabling end-to-end parameter tuning. Meanwhile, the dynamic of the information flow aligns with that of a standard stack. When $a_{t}^{\text{push}}$ dominates in $a_{t}$, all elements in the stack tend to shift downward as more information from $h_{t}$ flows into the top element; conversely, when the pop operation dominates, the elements in the stack shift upward as the information of the top element is mostly removed.
Stack mask
To implement our proposed hidden state stack using a list, the overall available stack size $S$ must be sufficiently large. Assuming $S$ is large enough, the tail of the stack is always padded with zero vectors. This padding ensures that stack operations comply with logical constraints and prevents invalid behaviors. However, this process is inherently discrete. To address this, we propose maintaining a differentiable stack mask for StackTrans. Specifically, the $i$-th element in the mask $\text{M}_{t}\in\mathbb{R}^{S}$ suggests how likely the corresponding element in the stack is active. The mask is updated with dynamics similar to those described in Equation 2:
$\text{M}_{t+1}[i]=\left\{\begin{array}[]{lclcll}a_{t}^{\text{push}}\cdot 1&+&a_{t}^{\text{pop}}\cdot\text{M}_{t}[1]&+&a_{t}^{\text{noop}}\cdot\text{M}_{t}[0],&\text{if }i=0\\[3.0pt] a_{t}^{\text{push}}\cdot\text{M}_{t}[i-1]&+&a_{t}^{\text{pop}}\cdot 0&+&a_{t}^{\text{noop}}\cdot\text{M}_{t}[i],&\text{if }i=S-1\\[3.0pt] a_{t}^{\text{push}}\cdot\text{M}_{t}[i-1]&+&a_{t}^{\text{pop}}\cdot\text{M}_{t}[i+1]&+&a_{t}^{\text{noop}}\cdot\text{M}_{t}[i],&\text{otherwise}\\[3.0pt] \end{array}\right.$ (3)
$\text{M}_{t}$ serves as an activation controller in StackTrans– if $a_{t}^{\text{push}}$ dominates, one more stack element is further activated, otherwise $a_{t}^{\text{pop}}$ dominates in $a_{t}$, the last activated element is more likely to be deactivated. When accessing the stack, we pad the stack by element-wise production of $\text{St}_{t}$ and $\text{M}_{t}$.
Global read
The standard stack simply peeks and returns the top element during reading. Although peeking is quite straightforward, we notice that it may cause unstable training during our initial experiments. Furthermore, limiting access to only the top element restricts gradient flow, reducing learning efficiency and leading to unstable training dynamics. A detailed discussion is provided in §B.4. Therefore, we propose the global read mechanism, by collecting information over the stack. Global read is achieved through a learnable query-over-stack attention:
$\text{R}_{t}=\text{Softmax}(W_{g}\cdot(\text{St}_{t}\otimes\text{M}_{t}))\cdot\text{St}_{t}$ (4)
where $\otimes$ refers to element-wise production and $W_{g}\in\mathbb{R}^{S}$ is a trainable query vector. The Softmax term computes the attention score, and the content read from the stack is the weighted sum of all stack elements, where each element is weighted by its corresponding attention score. The final output is a residual-like connection $h^{\prime}_{t}=g_{h}\cdot h_{t}+\text{R}_{t}$, where $g_{h}$ is a trainable parameter.
3.2 Multi-Head Low-Rank Stack
In the Transformer architecture, the multi-head attention mechanism [Vaswani et al., 2017] processes multiple attention patterns in parallel across different representation subspaces. The design enables the model to capture diverse relationships within the input sequence. Following a similar design philosophy, we propose the multi-head low-rank stack. Specifically, we down-project the hidden state $h_{t}\in\mathbb{R}^{d}$ and split it into subspaces ($\mathbb{R}^{d_{s}}$) as: $\left[h^{(1)}_{t},h^{(2)}_{t},\ldots,h^{(H)}_{t}\right]=\text{Reshape}(W_{\mathrm{down}}\cdot h_{t})$, where $H$ denotes the number of stack heads and $W_{\text{down}}\in\mathbb{R}^{(H\cdot d_{s})\times d}$ is the down-projection matrix. Each head corresponds to an independent stack as introduced in §3.1.
Given the down-projected hidden state $h_{t}^{(i)}\in\mathbb{R}^{d_{s}}$ for the $i$-th head, the stack element $\text{St}_{t}^{(i)}$ and the mask $\text{M}_{t}^{(i)}$ (both in $\mathbb{R}^{S\times d_{s}}$) along with the final read-out $\text{R}_{t}^{(i)}$ are computed independently for each head following Equations 1 - 4. After performing soft updates and global reads for all $H$ heads, we concatenate their outputs to obtain the final result: $\text{h}^{{}^{\prime}}_{t}=g_{h}\cdot h_{t}+W_{\mathrm{up}}\cdot\mathrm{Concat}\left(\text{R}^{(1)}_{t};\cdots;\text{R}^{(H)}_{t}\right)$, where $W_{\text{up}}\in\mathbb{R}^{d\times(H\cdot d_{s})}$ is the up-projection matrix. This multi-head mechanism allows the model to organize stack operations into different patterns in parallel, thereby capturing diverse relationships and dependencies more effectively. Empirically, we find that a small $H$ (e.g., 4 or 8) is sufficient to achieve notable performance. For more details on ablation studies and discussions, please refer to §6. On the other hand, the overall computational cost with the low-rank design is much smaller than counterpart of a single-head stack with full dimensions due to the reduced dimension of low-rank adaptation in each sub-stack. Refer to §6 for ablations.
3.3 Key Implementation & Training Know-Hows
The modular stack of StackTrans enables it to augment the Transformer architecture without altering the Transformer layers themselves. There are some key implementation and training insights.
Stack Overflow
In §3.1 and §3.2, we assume that the stack size $S$ is sufficiently large or even infinite. However, due to limitations in computational power and storage resources, $S$ is typically relatively small, making overflow inevitable. In our implementation, we address this by truncating the stack and setting all overflow elements to zero, that is, $\text{St}_{t}[i]=\vec{0}$ if $i\geq S$. This truncation can be seen as a form of “forgetting”, where the information carried by the overflow elements is discarded.
Training parallelism
Ideally, the stack is supposed to process hidden states according to their temporal or layer dependencies, i.e., it should prioritize hidden states generated from earlier tokens or shallower layers. One feasible sequence fed into the hidden state stack would be $\left[h_{{t_{0}},0},h_{{t_{0}},1},\cdots,h_{{t_{0}},L},h_{{t_{1}},0},\cdots\right]$, where $h_{t,i}\in\mathbb{R}^{d}$ denotes the hidden state of the $i$-th Transformer layer at token $t$, and $L$ represents the total number of layers. Although such a behavioral pattern is clearly beneficial for learning stack operations, the temporal dependencies conflict with the parallel training of the Transformer layers. To facilitate training parallelism, we implement StackTrans by breaking these temporal dependencies. Specifically, StackTrans learns stack operations based on the hidden state sequence at token $t_{i}$ from layer 00 to $L$ ($\left[h_{{t_{i}},0},\cdots,h_{{t_{i}},L}\right]$), allowing all tokens to be trained in parallel. We provide detailed discussion in §B.3.
Stack regularization
During training, StackTrans optimizes the standard autoregressive language modeling loss ($\mathcal{L}_{\text{LM}}$) over the token sequence. To prevent the operation probabilities $a_{t}$ from collapsing into uniform, we introduce an entropy-based regularization term, defined as $\mathcal{L}_{\text{St}}=\sum_{t}\mathcal{H}(a_{t})$, where $\mathcal{H}(\cdot)$ calculates the entropy. The overall loss function combines the language modeling loss and the stack regularization term, $\mathcal{L}=\mathcal{L}_{\text{LM}}+\lambda\cdot\mathcal{L}_{\text{St}}$, where $\lambda$ is a hyperparameter.
Table 1: Test accuracy on formal language tasks.
Task LSTM StackRNN StackAttn* Transformer StackTrans
Regular (RE) Tasks
Even Pairs 1.00 1.00 - 0.49 1.00
Parity Check 1.00 1.00 - 0.50 1.00
Cycle Navigation 0.89 1.00 - 0.20 1.00
Deterministic Context-Free (DCF) Tasks
Stack Manipulation 0.66 0.85 0.93 0.53 0.92
Reverse String 0.71 0.80 1.00 0.55 1.00
Modular Arithmetic 0.43 0.42 0.30 0.30 0.60
Context-Sensitive (CS) Tasks
Missing Duplicate 0.68 0.67 - 0.53 1.00
Odds First 0.60 0.55 - 0.51 0.53
Binary Addition 0.56 0.51 - 0.48 0.48
Binary Multiplication 0.56 0.53 - 0.50 0.48
Compute Sqrt 0.64 0.63 - 0.51 0.57
Bucket Sort 0.79 0.75 - 0.79 0.89
    Results listed here are reported by DuSell and Chiang [2024].
4 Evaluation against Formal Languages
Understanding formal languages is fundamental for modeling many aspects of real-world natural language processing tasks. To highlight the motivation behind our stack-enhanced mechanism, we first evaluate StackTrans on formal language modeling tasks inspired by the Chomsky hierarchy.
Experimental Setup
In this section, we evaluate StackTrans on three groups formal language modeling tasks aligned with the Chomsky hierarchy [Delétang et al., 2022]. These tasks assess a model’s ability to learn underlying compositional rules of formal languages and generalize to input lengths beyond those seen during training. Please refer to §D for details. Following prior work [Delétang et al., 2022], we implement StackTrans with relatively limited parameters. Concretely, we use five Transformer layers with $d=64$. $H$ is set to 4 and $d_{s}$ is set to 8. We compare StackTrans to some representative baselines, including the standard Transformer [Vaswani et al., 2017], LSTM [Graves and Graves, 2012], StackRNN [Joulin and Mikolov, 2015]) and StackAttn [Li et al., 2024], maintaining identical experimental settings for all models.
Evaluation Results
Table 1 shows that StackTrans consistently outperforms the standard Transformer, particularly on RE and DCF tasks. For RE tasks, most evaluated models attain near-perfect accuracy. However, Transformers tend to falter in the absence of explicit inductive biases. This further underscores the effectiveness of the hidden state mechanism introduced in StackTrans. When compared with the state-of-the-art stack-augmented approaches, such as StackRNN and StackAttn, StackTrans either outperforms them or is at least on par with them in nearly all tasks.
The hidden state stack mechanism likely endows StackTrans with characteristics akin to pushdown automata. This is evidenced by its superior performance over all baselines on both RE and DCF tasks, which are known to be solvable by pushdown automata. However, as pushdown automata are theoretically incapable of resolving CS tasks, we observe that all approaches, including StackTrans, perform poorly on CS tasks. Despite this limitation, StackTrans demonstrates the ability to handle a subset of CS tasks, which we attribute to specific design enhancements such as the multi-head mechanism and the global reading capability. These features provide StackTrans with stronger modeling capacity than traditional pushdown automata, enabling it to capture additional dependencies and complexities beyond the theoretical limits of pushdown automata.
5 Evaluation against General Natural Languages
From the perspective of computational linguistics, natural languages are generally considered to belong to a class of languages that includes DCF [Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985]. Therefore, to thoroughly assess StackTrans, we conduct further evaluations on general language modeling tasks. ❶ We examine the scalability of StackTrans through scaling law studies, analyzing the effect caused by model size and training tokens. ❷ A StackTrans with 360 million parameters is pretrained with nearly 980 billion tokens. We evaluate its performance on a variety of standard benchmarks, comparing it to models with similar or larger parameter scales. ❸ We provide additional empirical observations through ablation studies and deeper analysis (please refer to §6).
Experimental Setup
We follow the OLMo framework [allenai, 2024] to pretrain StackTrans. Our corpora comes from Dolma [Soldaini et al., 2024] and Smoll [Allal et al., 2025], which contain high-quality natural language, math, and Python code examples with diverse domains. We carry out data filtration, ultimately obtaining approximately 980 billion tokens. To scale up model parameters, we adapt the Dolma v1.6-sample configuration in OLMo, using roughly 80 billion tokens for each variant model training. For scaling up training tokens, we train StackTrans with 360M parameters on a sampled subset of 200 billion tokens from pretraining corpora.
Scaling Law of StackTrans
We train StackTrans models with a range of parameter sizes (360M, 600M, 1.0B, 1.5B, and 7B) under the same training budget in terms of tokens. The language modeling loss is tracked throughout the training process, and the scaling trends are depicted in Figure LABEL:fig:scalinglaw. Our observations find that StackTrans exhibits smoother convergence and attains lower final loss compared to standard Transformers of equivalent size. Notably, even with 360M parameters, StackTrans consistently demonstrates smaller loss, which underscores the significant contribution of the hidden state stack mechanism to improved generalization capabilities. Overall, StackTrans aligns well with the predicted scaling trends and delivers superior performance. To analyze the optimization process, we compare training loss dynamics between StackTrans and the standard Transformer in Figure LABEL:fig:loss. Detailed analysis is shown in §B.5.
Table 2: Evaluation results on natural language tasks.
StackTrans SmolLM SmolLM2 Qwen2.5 OLMo TinyLLaMA
Param. # (B) 0.36 0.36 0.36 0.5 1.1 1.1
Token # (T) 1 0.6 4 18 2 3
HellaSwag 59.5 51.8 54.5 51.2 60.7 55.2
ARC 59.0 50.1 53.0 45.4 44.0 43.4
PIQA 71.7 71.6 71.7 69.9 75.2 72.9
MMLU 36.5 34.4 35.8 33.7 31.9 32.2
CommonsenseQA 37.1 35.3 38.0 31.6 40.3 37.0
TriviaQA 11.2 9.1 16.9 4.3 2.8 9.8
Winogrande 52.8 52.8 52.5 54.1 53.2 55.7
OpenBookQA 37.5 37.2 37.4 37.4 38.0 33.2
GSM8K (5-shot) 33.6 1.6 3.2 33.4 1.8 1.7
Average 44.3 38.2 40.3 40.1 38.7 37.9
Evaluation against StackTrans-360M
We pre-train StackTrans-360M from scratch, and the detailed model configuration is shown in §F. To assess the downstream capabilities, we evaluate StackTrans-360M on a comprehensive suite of widely-used benchmarks, and details are shown in §E. As listed in Table 2, StackTrans-360M outperforms all baseline models, including those with significantly larger parameter sizes. Notably, it achieves substantial gains on GSM8K and ARC, highlighting its strength in reasoning tasks that require compositional generalization, recursion, or latent state management. Despite having fewer parameters, StackTrans performs competitively on PIQA and CommonsenseQA, further indicating that the stack-augmented memory module improves representation capacity without compromising generalization. Overall, StackTrans-360M achieves an average performance of 44.3 across 9 diverse tasks, exceeding comparable models in the table with only a fraction of the parameter size and dataset size. Our proposed StackTrans enhances LLM’s generalization ability, especially in scenarios with limited computation and parameter budgets.
6 Discussion
Ablations of Key Designs
To assess the impact of the stack design in StackTrans, we investigate two alternative configurations. ❶ We replace the stack with a queue, adopting a first-in-first-out storage structure, which we term QueueTrans. Apart from this modification, the queue mask performs similar functions to the stack mask in maintaining valid and activated elements. ❷ In this extreme setting, we modify the stack in StackTrans to “push-only”, where $a_{t}^{push}$ is fixed to 1. This configuration essentially disables pop and no-op operations. The multi-head stack and the low-rank compression are the other two crucial design elements in StackTrans. To study their impact, we conduct ablation studies as follows: ❸ We disable the multi-head splitting, reverting it back to a single-head stack. ❹ We remove the down- and up-projections (low-rank mechanism) from the full stack. In total, we create four variants for ablation studies.
Table 3: Training and validation results of stack ablations ($\sim$20B tokens).
Model Train Loss $\downarrow$ V2 Loss $\downarrow$ V2 PPL $\downarrow$ V3 Loss $\downarrow$ V3 PPL $\downarrow$
Transformer 2.411 3.518 34.38 3.195 25.33
StackTrans 2.359 3.432 32.89 3.092 24.50
[…]
Push-Only (Fix $a_{t}^{\text{push}}=1$) 2.875 4.032 39.02 3.407 27.13
Single-Head ($H=1$) 2.493 3.552 33.56 3.130 24.91
Full-Dimension ($H\cdot d_{s}=d$) 2.370 3.457 33.05 3.105 24.73
We evaluate all the variants introduced above on the V2 and V3 validation sets [Zhu et al., 2024]. Experimental details are provided in §G. The ablation results are presented in Table 3. The QueueTrans variant exhibits notably higher perplexity and lower overall accuracy compared to StackTrans, particularly on tasks involving hierarchical or recursive patterns. This outcome is consistent with our expectation that queue operations are inherently less effective at modeling nested dependencies and grammars. Similarly, the push-only variant performs poorly, with both training and validation losses significantly deteriorating. The absence of pop operations impairs its ability to dynamically manage and retrieve stored information, thereby reducing its overall effectiveness. Both single-head and full-dimensional variants are consistently outperformed by StackTrans. The multi-head mechanism enhances flexibility by enabling parallel decomposition of stack streams, while the low-rank mechanism reduces computational costs without compromising much modeling capacity.
Ablations of Hyperparametesr
In addition to the configuration of the standard Transformer layers, StackTrans has three key hyperparameters – the number of stack heads $H$, the dimension of each stack head $d_{s}$, and the stack size $S$. We perform grid search over reasonable ranges for these hyperparameters, training StackTrans with 20 billion tokens. The curves of training loss and validation loss (evaluated on V2 [Zhu et al., 2024]) are plotted in Figure LABEL:fig:stackablation. It is clear that $H$ is crucial for parallelism, but Figure LABEL:fig:ablation:head indicates that performance plateaus once $H$ surpasses a certain threshold. From Figure LABEL:fig:ablation:dim, we observe that setting $d_{s}$ within the range from 16 to 64 balances the computational cost and the model’s expressiveness effectively. Similarly, Figure LABEL:fig:ablation:slots shows that increasing $S$ from 24 to 32 has nearly no impact on performance. Given that a larger $S$ leads to higher storage overhead, we make a trade-off and ultimately set $S$ to 24 during our evaluation.
More Detailed Discussion
Due to the length constraints of the paper, we provide more discussions in the appendices. For in-depth investigations of StackTrans’s training and inference efficiency, please refer to §B.1. In our implementation, we break the temporal dependencies to facilitate training parallelism of StackTrans, as briefly introduced in §3.3. We further discuss why this approximation works in §B.3. We adopt global reading rather than top peeking for StackTrans, and we explain the rationale behind this design in §B.4. We provide visualizations of the stack action patterns across different tasks in §B.2, and analyze the training dynamics in §B.5.
7 Conclusion
Inspired by pushdown automata, we propose StackTrans, a novel Transformer variant architecture integrating differentiable hidden state stacks in between Transformer layers. StackTrans improves generalization in both formal language tasks and natural language modeling tasks. In particular, our from-scratch pretrained StackTrans-360M outperforms several larger open-source LLMs with 2–3$\times$ more parameters, showcasing its superior efficiency and reasoning capability.
Appendix A Related Work
Evolvement of Large Language Models
The success of LLMs is deeply rooted in the Transformer architecture [Vaswani et al., 2017] and its subsequent variations, which serve as the backbone for LLMs such as GPT [Radford et al., 2018; GPT-4, 2023; GPT-4o, 2024], DeepSeek [Bi et al., 2024; Liu et al., 2024], LLaMA [Touvron et al., 2023], and other prominent LLM series [Brown et al., 2020]. Scaling laws [Kaplan et al., 2020] have shown that increasing model parameter size leads to emergent capabilities, enabling LLMs to tackle increasingly complex tasks and exhibit surprising generalization behaviors. In addition to proprietary closed systems, open-source initiatives like OLMo [Groeneveld et al., 2024] demonstrate the potential of community-scale pretraining using meticulously curated datasets such as Dolma [Soldaini et al., 2024]. These efforts highlight how transparent methodologies can accelerate innovation, making powerful LLMs more accessible for academic research and industrial applications. Despite their widespread adoption, the Transformer architecture has inherent expressivity limitations [Hahn, 2020]. Although Transformers are theoretically Turing complete [Chomsky, 1956], they often underperform on tasks tied to formal languages within the Chomsky hierarchy. Recent studies [Delétang et al., 2022; DuSell and Chiang, 2024] have shown that standard Transformers struggle with recursive and hierarchical patterns across both synthetic and real-world tasks. These limitations underscore the need for fundamental architectural enhancements to better model the Chomsky hierarchy, such as rich linguistic and algorithmic structures.
Building on prior work and drawing inspiration from pushdown automata [DuSell and Chiang, 2024; Li et al., 2024; Sartran et al., 2022], we address these limitations by introducing hidden state stacks into the Transformer architecture. Our proposed method, StackTrans, enhances the model capacity to represent hierarchical dependencies and recursive grammars, enabling it to learn Chomsky hierarchy grammars effectively. We believe that fostering transparent and open discussions around the underlying architectural challenges will accelerate the evolution of Transformer-based models and propel the development of large language models to new heights.
Stack Augmentation
Equipping neural networks with external data structures, such as stacks, has been widely explored to enhance models’ ability to recognize hierarchical and context-free languages. Although earlier studies primarily focused on recurrent neural networks [Joulin and Mikolov, 2015; Grefenstette et al., 2015], recent efforts have adapted these thoughts to Transformer-based models [DuSell and Chiang, 2024; Li et al., 2024; Sartran et al., 2022]. They aim to embed stack-like operations into Transformers to address their shortcomings in modeling the Chomsky hierarchy, particularly DCFs. For example, Li et al. [2024] augment standard attention layers with differentiable stacks, enabling soft push/pop operations to model recursive structures. However, this comes with architectural trade-offs, where the stack control is tightly coupled with the attention mechanism, leading to increased entanglement and reduced modularity. Further advances, such as those in DuSell and Chiang [2024], embed stack operations directly within attention heads, providing stronger inductive biases. While these approaches better model Chomsky hierarchy grammars, their validation is largely limited to small-scale models and synthetic datasets. It raises questions about their scalability and generalization ability to large-scale tasks and larger Transformer-based models with millions or billions of parameters. In contrast, StackTrans introduces a differentiable stack. Rather than tempering the attention mechanism, StackTrans integrates stacks in between standard Transformer layers. This design decouples stack manipulation from attention computations, enabling seamless integration with pretrained Transformer models while preserving compatibility across architectures of varying sizes (ranging from 0.36B to 7B parameters). By focusing on hidden states, StackTrans maintains flexibility and scalability, addressing broader linguistic and algorithmic challenges without compromising the core principles of the Transformer architecture.
Appendix B Discussion (Cont.)
Following §6, we provide some more discussions in this section.
B.1 Training and Inference Efficiency
Table 4: Training and inference efficiency, including time cost and GPU memory usage.
Model Variant Time Cost Peak GPU Memory Usage
Training Inference
Transformer* $\times$1.00 $\times$1.00 $\times$1.00
StackTrans $\times$1.16 $\times$1.09 $\times$1.12
Single-Head Stack $\times$1.03 $\times$1.04 $\times$1.00
Full-Dimensional Stack $\times$3.78 $\times$3.30 $\times$1.73
    All values represent multiples of results of the baseline Transformer. Constrained by further engineering implementation (which we are actively working on), we conducted evaluations without enabling various cache or optimization mechanisms.
Considering that the training and inference efficiency is an important factor for model applications. In this section, we investigate the training and inference efficiency of StackTrans compared to the standard Transformer. To keep a fair comparison, all comparisons are conducted on the same hardware setup. We measure both training and inference time over 100 consecutive steps under identical hyperparameter and batch size configurations. To explore the design trade-offs, we also compare several StackTrans variants, including single-head stack and full-dimensional stack as described in §6. As shown in Table 4, despite incorporating differentiable stack modules, StackTrans achieves competitive training and inference efficiency. Concretely, it introduces only marginal overhead (around 10%) compared to the standard Transformer while yielding significant performance improvements. Besides, the memory usage increase is moderate, well within the typical consumption range of large-scale LLM. This suggests that StackTrans offers a practical and scalable approach, and its stack mechanism can be integrated into existing Transformer without compromising deployment efficiency.
B.2 Stack Action Patterns across Layers and Tasks
Figure 9: Average probabilities of three operations across the network depth for the two-digit addition task and the two-digit multiplication task.
StackTrans introduces layer-wise stack modules that manipulate memory using three soft actions: push, pop, and no-op. Since stack dynamics play a central role in the model’s expressiveness, we conduct an in-depth analysis of the stack action patterns across layers and downstream tasks. We select two arithmetic tasks from our synthetic benchmark suite: the two-digit addition task and the two-digit multiplication task. Although both tasks involve structured numerical reasoning, multiplication generally requires deeper or more nested intermediate steps than addition. Our StackTrans-360M achieves 100% accuracy on both tasks, indicating the performance of our model on these basic arithmetic questions. For every layer and timestep, we compute the average action probabilities of the three operations for each Transformer layer in StackTrans and visualize their trends across the network depth, as shown in Figure 9.
The results reveal a consistent action distribution pattern on both tasks: earlier layers predominantly favor push operations, while later layers exhibit an increased use of pop, with no-op remaining relatively stable throughout. This trend suggests that StackTrans automatically learns to incrementally store information during lower layers and retrieve it in upper layers. The intuitive behavior mirrors how hierarchical or recursive structures are processed.
Figure 9 further shows that multiplication elicits markedly more push operations in middle layers and deferred pop activity in higher layers, reflecting the deeper computation graph required by the task. In contrast, addition induces a flatter push/pop pattern distributed more evenly across layers, consistent with its shallower reasoning structure. These findings confirm that StackTrans learns to adapt memory access patterns dynamically according to task complexity, and that its stack behavior is both interpretable and task-sensitive.
B.3 Approximations for Training Parallelism
In our implementation, we introduce necessary approximations to maintain training parallelism while preserving model performance, as detailed in §3.3. Temporal dependencies among hidden states are a fundamental aspect of the Transformer’s processing pipeline. Let $\mathbf{h}_{t,i}\in\mathbb{R}^{d}$ represent the hidden state at layer $i$ for token $t$, where the ideal processing order for our stack would follow the complete sequence:
$[\mathbf{h}_{t_{0},0},\mathbf{h}_{t_{0},1},\ldots,\mathbf{h}_{t_{0},L},\mathbf{h}_{t_{1},0},\ldots],$ (5)
with $L$ denoting the total number of layers. However, to enable practical training parallelism, we introduce a controlled truncation between $\mathbf{h}_{t,L}$ and $\mathbf{h}_{t+1,0}$. This approximation allows us to compute token losses for all elements in a sequence simultaneously, which would otherwise be computationally prohibitive.
The decision to break these temporal dependencies is guided by two key considerations. First, the self-attention mechanism inherently captures cross-token relationships, which can partially compensate for the truncation. Second, the stack mechanism introduced in our model complements the attention layers by retaining sequential dependencies through external memory operations. Empirical evidence from prior work [Kudo et al., 2024] supports this design, showing that those intermediate-layer hidden states for subsequent tokens effectively preserve information from earlier tokens. Overall, this approximation achieves an optimal trade-off between computational efficiency and model performance, enabling scalable and parallelizable training while maintaining our designed stack mechanism.
B.4 Global Reading Capability
Traditional stacks only permit access to the top element. However, in neural network modeling, such a restriction is unnecessary since tensor vectors can be efficiently accessed through operations like matrix multiplication. To enhance the stack’s representational ability, our differential stack eliminates this constraint by introducing global reading capabilities and enabling full random access.
Furthermore, we find that enforcing a strict top-only access during training leads to unstable and suboptimal model performance. We attribute this to the frequent stack operations in neural networks: limiting access to the top element disrupts gradient flow and reduces parameter learning efficiency. By relaxing this constraint and enabling global read operations, our differential stack achieves greater representational ability and improved adaptability across diverse tasks.
B.5 Training Loss Dynamics
To analyze the optimization process, we conduct a comparative analysis of training loss dynamics between StackTrans and the standard Transformer. The training curves are presented in Figure LABEL:fig:loss. One may find out that StackTrans exhibits a slightly slower decrease in loss during the very early training stages compared to the standard Transformer. We attribute this behavior to the additional learning complexity introduced by the hidden state stack, where StackTrans must learn when and how to carry out stack operations. As training progresses, however, StackTrans not only catches up but eventually surpasses the standard Transformer in convergence speed. Once the stack operation distribution stabilizes, StackTrans begins to leverage the stack more effectively, leading to a steeper decline in loss and an overall lower convergence plateau. This phenomenon shows that StackTrans, while requiring a slightly longer warm-up phase, ultimately achieves greater learning efficiency and a superior asymptotic performance ceiling compared to the standard Transformer.
Appendix C Limitations
While StackTrans demonstrates strong performance across a variety of tasks, there are several limitations to our work that we aim to address.
Table 5: Formal language task descriptions and input-output examples.
Task Name Description
Regular (RE) Tasks
Even Pairs Check if the count of ab/ba pairs is even.
Parity Check Check if the count of b is even.
Cycle Navigation Navigate movements on a modulo-5 cycle.
Deterministic Context-Free (DCF) Tasks
Stack Manipulation Perform stack operations and return the final state.
Reverse String Reverse the input string using a stack.
Modular Arithmetic Evaluate nested arithmetic expressions modulo 5.
Solve Equation Find a variable satisfying a modular equation.
Context-Sensitive (CS) Tasks
Binary Addition Compute binary addition of two numbers.
Binary Multiplication Compute binary multiplication.
Compute Sqrt Compute the integer square root of a binary number.
Bucket Sort Sort a sequence over a fixed alphabet.
Duplicate String Output the string concatenated with itself.
Missing Duplicate Find the missing character in a duplicated string.
Odds First Interleave odd and even indices of a sequence.
Task Name Input Example Output Example
Regular (RE) Tasks
Even Pairs aabba True
Parity Check aaabba True
Cycle Navigation 011210 2
Deterministic Context-Free (DCF) Tasks
Stack Manipulation abbaa POP PUSH a POP abba
Reverse String aabba abbaa
Modular Arithmetic $-(1-2)\cdot(4-3\cdot(-2))$ 0
Solve Equation $-(z-2)\cdot(4-3\cdot(-2))=0$ 1
Context-Sensitive (CS) Tasks
Binary Addition 10010 + 101 10111
Binary Multiplication 10010 $\times$ 101 1001000
Compute Sqrt 101001 101
Bucket Sort 421302214 011222344
Duplicate String abaab abaababaab
Missing Duplicate ab_aba a
Odds First aaabaa aaaaba
Limitations of Model Size and Dataset Size
Constrained by computational resources, we limit our final pre-trained model to 360M parameters and use approximately 1 trillion training tokens. Although the results show competitive performance, the scaling law discussed in §5 suggests that larger models and datasets could further amplify the strengths of StackTrans. Particularly, scaling up the number of model parameters and training tokens may enhance its ability to tackle more complex tasks. This limitation highlights the importance of access to large-scale computing infrastructure for future research. We hope to leverage the power of the open-source community to validate this new architecture.
Necessary Approximation in Design
To achieve training parallelism, we introduce controlled approximations in the sequence processing pipeline, as detailed in §3.3. Specifically, the truncation of temporal dependencies between tokens facilitates scalable training but may reduce the model’s ability to fully exploit fine-grained sequential patterns. While the self-attention mechanism and stack-based memory mitigate this limitation, the truncation approximation may still pose risks, especially in tasks requiring deep inter-token dependencies. We plan to explore more robust and efficient settings in future work.
Appendix D Formal Language Task Details
We follow the experimental settings of Delétang et al. [2022]. Table 5 presents an overview of the formal language tasks and their complexity level within the Chomsky hierarchy. These tasks assess models’ ability to learn underlying compositional rules of formal languages and generalize to input lengths beyond those seen during training. The three task groups are categorized by the Chomsky hierarchy type, including RE (type-3), DCF (type-2), and CS (type-1). The classification adheres to formal automata theory, associating the three tasks with finite-state automata, pushdown automata, and linear-bounded automata, respectively. Please refer to Table 5 in §D for definitions and examples of these tasks. Despite the presence of classification tasks, all tasks are formulated as sequence mapping problems. In this setup, the model takes an input sequence and decodes it into an output sequence. StackTrans is trained on sequences with input length uniformly sampled from 1 to 40 tokens. At test time, we evaluate StackTrans on sequences with significantly longer lengths up to 500 tokens, thereby measuring its length generalization. Following the same procedure as Delétang et al. [2022], token-level accuracy is used as the evaluation metric. We repeat each experimental configuration ten times and report the best accuracy achieved.
Table 6: Model configuration of StackTrans-360M.
Parameter Value
Vocabulary Size 49152
Number of Attention Heads 15
Number of Hidden Layers 32
Hidden Size 960
Intermediate Size (FFN) 2560
Attention Dropout 0.0
Activation Function Silu
Number of Stack Heads 4
Stack Dimensionality 16
Stack Size 24
Maximum Position Embeddings 4096
RoPE Scaling None
RoPE $\theta$ 100000
Table 7: Overview of V2 and V3 Validation Sets. Each validation set includes diverse text sources to ensure comprehensive evaluation.
Validation Set Datasets Included
V2 Validation Sets
v2-small-4chan-validation, v2-small-c4_100_domains-validation,
v2-small-c4_en-validation, v2-small-gab-validation,
v2-small-ice-validation, v2-small-m2d2_s2orc-validation,
v2-small-m2d2_wiki-validation, v2-small-manosphere-validation,
v2-small-mc4_en-validation, v2-small-pile-validation,
v2-small-ptb-validation, v2-small-twitterAEE-validation,
v2-small-wikitext_103-validation
V3 Validation Sets
v3-small-c4_en-validation, v3-small-dolma_books-validation,
v3-small-dolma_common_crawl-validation,
v3-small-dolma_pes2o-validation, v3-small-dolma_reddit-validation,
v3-small-dolma_stack-validation, v3-small-dolma_wiki-validation,
v3-small-ice-validation, v3-small-m2d2_s2orc-validation,
v3-small-pile-validation, v3-small-wikitext_103-validation
Appendix E General Natural Language Task Details
To assess the downstream capabilities, we evaluate StackTrans-360M on a comprehensive suite of widely-used benchmarks [Brown et al., 2020; Touvron et al., 2023; Groeneveld et al., 2024], including those for common sense reasoning (e.g., HellaSwag, PIQA) [Zellers et al., 2019; Clark et al., 2018; Hendrycks et al., 2021; Bisk et al., 2020; Sakaguchi et al., 2020], question answering (e.g., OpenBookQA) [Talmor et al., 2019; Joshi et al., 2017; Mihaylov et al., 2018], and math-based reasoning (GSM8K) [Cobbe et al., 2021]  [^2] The evaluation protocol strictly follows Allal et al. [2025], and we obtain nearly identical results to those reported in the paper.. We compare StackTrans-360M with other open-source models around 1B parameters, including SmolLM-360M [Allal et al., 2024], SmolLM2-360M [Allal et al., 2025], Qwen2.5-0.5B [Yang et al., 2024], OLMo-1B [Groeneveld et al., 2024], and TinyLLaMA-1B [Zhang et al., 2024b]. We use the lighteval framework [huggingface, 2024], and for all applicable tasks, we adhere to zero-shot evaluation settings, unless otherwise specified.
Appendix F Model Configuration of StackTrans-360M
Table 6 shows the detailed model configuration of our StackTrans-360M, which is inspired by the similar setting in Allal et al. [2024] and Allal et al. [2025]. The stack-related setting is decided by a grid search in §6.
Appendix G Validation Dataset Details for General Language Modeling
Following the method of Zhu et al. [2024], we evaluate all variants on the V2 Validation Sets and V3 Validation Sets curated within the OLMO framework. The specific datasets for V2 and V3 validation [Zhu et al., 2024] are shown in Table 7.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claim for this paper is to enhance LLMs with the stack mechanism. The content of this paper focuses on the design and evaluation of this method, which is the main contribution of this paper. We give a full contribution list in §3(a).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in §C.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The theoretical result and proof is mainly from existing related work in §A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed experimental setup for each task in §4, §5, §D and §F.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will provide the key code for dataset processing and training in the supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all the training and test details for each task in §4, §5, §D and §F.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: All the evaluation settings strictly follow the default settings, and details are provided in §4 and §5. The randomness in these evaluations is small.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide all the model size and dataset setting for each task in §4, §5, §D and §F.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The dataset we use follows Apache-2.0 license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: This paper is an enhancement method for LLMs. The LLM is the basis of our method.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.